please answer this :write an equation of the line that passes through the given two points.
(2,3),(8,0)

Answer by venugopalramana(3286)

please answer this :write an equation of the line that passes through the given two points.
P(X1,Y1)=(2,3),Q=(X2,Y2)=(8,0)
EQN. IS GIVEN BY FORMULA
(Y-Y1)/(Y2-Y1)=(X-X1)/(X2-X1)
(Y-3)/(0-3)=(X-2)/(8-2)
Y-3=-3(X-2)/6=-(X-2)/2
2(Y-3)=-(X-2)
2Y-6+X-2=0
2Y+X-8=0